OpenStudy (gracia):

If y= sin2 2x, find dy/dx

OpenStudy (anonymous):

two of many options:\[\sin^2 (2x) = (\sin(2x))^2\] or \[\sin^2(2x) = \frac{ 1 - \cos(4x) }{ 2 }\]

OpenStudy (gracia):

thank you.

OpenStudy (gracia):

Find the first derivative:Y=sin(4x-5)

OpenStudy (anonymous):

4cos(4x-5)

OpenStudy (cggurumanjunath):

y'=2*2sin^2(2x)=4sin^2(2x)

OpenStudy (gracia):

Evaluate the limit: lim(-15) x 3

OpenStudy (gracia):

What is the minium value of thefunction F defined by F(x)=x2-6x+8

OpenStudy (gracia):

Help me...How high is a semi-elliptical arch of height 20ft and span 80 ft. at a distamce of 15 ft from the end.